OpenStudy (abbycross167):

Can someone please check my work? Please? I would really appreciate it... Question~ Solve the systems using the substitution method. x - y = 0 3x + y = -4

OpenStudy (abbycross167):

My answer~ Solve one equation for either variable~ Y=0+x In the other equation replace the variable from step 1 with what is equaled: 3x+(0+x)=-4 Solve the equation in step 2 for the other variable: 4x= -4 --- x= 4/-4 = -1 Substitute the value found in step 3 into equation in step 1: Y=0+-1 Solve equation in step 4: Y=-1 Write your answer as an ordered pair (x, y). (4/-4, 0/-1)

OpenStudy (abbycross167):

What my teacher said was wrong~ when you move the x in the 1st equation, it isn't +x

OpenStudy (sparrow2):

why you plug y ase 0/-1 when it is -1?

OpenStudy (mathmale):

Please reduce your answer. You end up with the coordinates of a point. Check these coordinates in the original equation. Is the resulting equation true or false? What I'd do is actually similar to what you've already done. Solving the first equation for y results in y=x. Substituting this into the second equation results in 3x + (x) = -4. Abby, please combine the 2 terms on the left. What equation do you obtain?

OpenStudy (mathmale):

No need to insert " 0 " into either of these equations.

OpenStudy (mathmale):

Your "In the other equation replace the variable from step 1 with what is equaled: 3x+(0+x)=-4" is fine; your inserting that ' 0 ' doesn't hurt, but doesn't help either. Toss out the (0+x) and replace it with just " x "

OpenStudy (abbycross167):

@mathmale I changed the 3x+(0+x)=-4 to 3x+x=-4, is that all I have to fix?